# PAC-Bayesian Adversarially Robust Generalization Bounds for Graph Neural Network

### v1

@@ A.4.1 Proof of Perturbation Bound for MPGNN
Adding perturbation $\Delta w$ to the weights $w$, i.e., for the $j$-th layer, the perturbed weights are $W_{j}+\Delta W_{j}$ or $U_{j}+\Delta U_{j}$ and denote $H_{j}^{\prime},\ E_{j}^{\prime},\ F_{j}^{\prime}$ as the perturbed term respectively. Upper Bound on the Node Representation. For any $j<l$,
@@
$\left\|f_{w+u}\left(G\right)-f_{w}\left(G\right)\right\|_{2}=\left\|H_{l}^{\prime}-H_{l}\right\|_{2}$$\leq\begin{cases}e\eta LM_{1}M_{2}\left\|X\right\|_{2}(l+1)^{2},&\tau=1,\\ e\eta lLM_{1}M_{2}\left\|X\right\|_{2}\dfrac{\tau^{l-1}-1}{\tau-1},&\tau\neq 1.\\ \end{cases}\qed$
@@
$\left|RM\left(f_{\tilde{w}}\left(G\right),i,j\right)-RM\left(f_{w}\left(G\right),i,j\right)\right|$ $=\left|\inf_{G^{\prime}\in\delta_{\tilde{w}}(G)}M\left(f_{\tilde{w}}\left(G\right),i,j\right)-\inf_{G^{\prime}\in\delta_{w}(G)}M\left(f_{w}\left(G\right),i,j\right)\right|$
@@
A.4.2 Proof of Generalization bounds for MPGNN
@@
We consider the cases of different values of $\tau$. Case $\tau\neq 1$ Let $\beta=\max\left\{\frac{1}{\zeta},\xi^{1/l}\right\}$, where $\zeta=\min\left(\left\{\left\|U_{i}\right\|_{2}\right\}_{i=1}^{l-1},\left\{\left\|W_{i}\right\|_{2}\right\}_{i=1}^{l}\right)$ and $\xi=LM_{1}M_{2}\dfrac{\tau^{l-1}-1}{\tau-1}$. Consider the prior distribution $P=\mathcal{N}(0,\sigma^{2}I)$ and the random perturbation distribution $u\sim\mathcal{N}(0,\sigma^{2}I)$. Note that the $\sigma$ of the prior and the perturbation distributions are the same and will be set according to $\beta$. More precisely, we will set the $\sigma$ based on some approximation $\tilde{\beta}$ of $\beta$ since the prior distribution $P$ can not depend on any learned weights directly. The approximation $\tilde{\beta}$ is chosen to be a cover set which covers the meaningful range of $\beta$. For now, let us assume that we have a fix $\tilde{\beta}$ and consider $\beta$ which satisfies $|\beta-\tilde{\beta}|\leq\frac{1}{l+2}\beta$. Note that this also implies
@@
$\mathbb{P}\left(\bigcap_{i=1}^{l-1}\left\{\left\|\Delta U_{i}\right\|_{2}<t\right\}\cap\bigcap_{i=1}^{l}\left\{\left\|\Delta W_{i}\right\|_{2}<t\right\}\right)$ $=1-\mathbb{P}\left(\bigcup_{i=1}^{l-1}\left\{\left\|\Delta U_{i}\right\|_{2}\geq t\right\}\cup\bigcup_{i=1}^{l}\left\{\left\|\Delta W_{i}\right\|_{2}\geq t\right\}\right)$
@@
and one of the sufficient conditions for this is $\frac{\gamma}{4e(B+\epsilon)}\leq\beta^{l}$. We will see how to satisfy this condition later. We now compute the KL term in the PAC-Bayes bound in Lemma 3.2. As defined above, for a fixed $w$, the prior distribution $P=\mathcal{N}(0,\sigma^{2}I)$, the posterior distribution $Q$ is the distribution of $\Delta w$ shifted by $w$, that is, $Q=\mathcal{N}(w,\sigma^{2}I)$. Then we have,
@@
[…]
@@
where $\hat{G}=\arg\inf_{G^{\prime}\in\delta_{w}(G)}M\left(f_{w}\left(G^{\prime}\right),i,j\right)$ and $H_{l-1}^{*}$ is the corresponding node representation. Therefore, by the definition, we always have $\hat{L}_{S,\gamma}\left(f_{w}\right)=1$ when $\beta^{l}<\frac{\gamma}{2(B+\epsilon)}$. Alternatively, if $\beta^{l}>\frac{\gamma\sqrt{m}}{2(B+\epsilon)}$, the term inside the big-O notation in Equation 18 would be
@@
And note that the lower bound in Equation 19 ensures that Equation 17 holds which in turn justifies the applicability of Lemma A.6. Consider a covering set centered at $C=\left\{\tilde{\beta}_{1},\dots,\tilde{\beta}_{|C|}\right\}$ of the interval in Equation 19 with radius $\frac{1}{l+2}\left(\frac{\gamma}{2(B+\epsilon)}\right)^{1/l}$, where $|C|$ is the size of the set. To make sure that $|\beta-\tilde{\beta}|\leq\frac{1}{l+2}\left(\frac{\gamma}{2(B+\epsilon)}\right)^{1/l}\leq\frac{1}{l+2}\beta$, the minimal size of $C$ should be
@@
$\leq\hat{R}_{S,\gamma}\left(f_{w}\right)+\mathcal{O}\left(\sqrt{\frac{(B+\epsilon)^{2}\max\left\{\zeta^{-1},\xi^{1/l}\right\}^{2l+2}l^{2}h\log(lh)\left\|w\right\|_{2}^{2}+\log(m(l+2)/\delta^{\prime})}{\gamma^{2}m}}\right),$
@@
and one of the sufficient conditions is $\frac{l\gamma}{4e(B+\epsilon)(l+1)^{2}}\leq\beta^{2}$. We will see how to satisfy this condition later. We now compute the KL term in the PAC-Bayes bound in Lemma 3.2. As defined above, for a fixed $w$, the prior distribution $P=\mathcal{N}(0,\sigma^{2}I)$, the posterior distribution $Q$ is the distribution of $\Delta w$ shifted by $w$, that is, $Q=\mathcal{N}(w,\sigma^{2}I)$. Then we have,
@@
where $\hat{G}=\arg\inf_{G^{\prime}\in\delta_{w}(G)}M\left(f_{w}\left(G^{\prime}\right),i,j\right)$ and $H_{l-1}^{*}$ is the corresponding node representation. Therefore, by the definition, we always have $\hat{R}_{S,\gamma}\left(f_{w}\right)=1$ when $\beta^{2}<\frac{\gamma}{2(B+\epsilon)l}$. Alternatively, if $\beta^{2}>\frac{\gamma\sqrt{m}}{2(B+\epsilon)l}$, the term inside the big-O notation in Equation 21 would be
@@
And note that the lower bound in Equation 22 ensures that Equation 20 holds which in turn justifies the applicability of Lemma A.6. Consider a covering set $C=\left\{\tilde{\beta}_{1},\dots,\tilde{\beta}_{|C|}\right\}$ of the interval in Equation 22 with radius $\frac{1}{4}\sqrt{\frac{\gamma}{2(B+\epsilon)l}}$, where $|C|$ is the size of the set. To make sure that $|\beta-\tilde{\beta}|\leq\frac{1}{4}\sqrt{\frac{\gamma}{2(B+\epsilon)l}}\leq\frac{1}{4}\beta$, the minimal size of $C$ should be
@@
$\leq\hat{L}_{S,\gamma}\left(f_{w}\right)+\mathcal{O}\left(\sqrt{\frac{(B+\epsilon)^{2}\max\left\{\zeta^{-6},\left(LM_{1}M_{2}\right)^{3}\right\}(l+1)^{4}h\log(lh)\left\|w\right\|_{2}^{2}+\log(m/\delta)}{\gamma^{2}m}}\right).\qed$

### v2

@@ A.3.1 Proof of Perturbation Bound for MPGNN
Adding perturbation $\Delta w$ to the weights $w$, i.e., for the $j$-th layer, the perturbed weights are $W_{j}+\Delta W_{j}$ or $U_{j}+\Delta U_{j}$ and denote $H_{j}^{\prime},\ E_{j}^{\prime},\ F_{j}^{\prime}$ as the perturbed term respectively.
Upper Bound on the Node Representation. For any $j<l$,
@@
$\left\|f_{w+u}\left(G\right)-f_{w}\left(G\right)\right\|_{2}=\left\|H_{l}^{\prime}-H_{l}\right\|_{2}$$\leq\begin{cases}e\eta LM_{1}M_{2}\left\|X\right\|_{2}(l+1)^{2},&\tau=1,\\ e\eta lLM_{1}M_{2}\left\|X\right\|_{2}\dfrac{\tau^{l-1}-1}{\tau-1},&\tau\neq 1.\\ \end{cases}$
@@
$\quad\,\left|RM\left(f_{\tilde{w}}\left(G\right),i,j\right)-RM\left(f_{w}\left(G\right),i,j\right)\right|$
$=\left|\inf_{G^{\prime}\in\delta_{\tilde{w}}(G)}M\left(f_{\tilde{w}}\left(G\right),i,j\right)-\inf_{G^{\prime}\in\delta_{w}(G)}M\left(f_{w}\left(G\right),i,j\right)\right|$
@@
A.3.2 Proof of Generalization bounds for MPGNN
@@
We consider the cases of different values of $\tau$.
Case $\tau\neq 1$ Let $\beta=\max\left\{\frac{1}{\zeta},\xi^{1/l}\right\}$, where $\zeta=\min\left(\left\{\left\|U_{i}\right\|_{2}\right\}_{i=1}^{l-1},\left\{\left\|W_{i}\right\|_{2}\right\}_{i=1}^{l}\right)$ and $\xi=LM_{1}M_{2}\dfrac{\tau^{l-1}-1}{\tau-1}$. Consider the prior distribution $P=\mathcal{N}(0,\sigma^{2}I)$ and the random perturbation distribution $u\sim\mathcal{N}(0,\sigma^{2}I)$. Note that the $\sigma$ of the prior and the perturbation distributions are the same and will be set according to $\beta$. More precisely, we will set the $\sigma$ based on some approximation $\tilde{\beta}$ of $\beta$ since the prior distribution $P$ cannot depend on any learned weights directly. The approximation $\tilde{\beta}$ is chosen to be a cover set which covers the meaningful range of $\beta$. For now, let us assume that we have a fix $\tilde{\beta}$ and consider $\beta$ which satisfies $|\beta-\tilde{\beta}|\leq\frac{1}{l+2}\beta$. Note that this also implies
@@
$\quad\,\mathbb{P}\left(\bigcap_{i=1}^{l-1}\left\{\left\|\Delta U_{i}\right\|_{2}<t\right\}\cap\bigcap_{i=1}^{l}\left\{\left\|\Delta W_{i}\right\|_{2}<t\right\}\right)$
$=1-\mathbb{P}\left(\bigcup_{i=1}^{l-1}\left\{\left\|\Delta U_{i}\right\|_{2}\geq t\right\}\cup\bigcup_{i=1}^{l}\left\{\left\|\Delta W_{i}\right\|_{2}\geq t\right\}\right)$
@@
and one of the sufficient conditions for this is $\frac{\gamma}{4e(B+\epsilon)}\leq\beta^{l}$. We will see how to satisfy this condition later.
We now compute the KL term in the PAC-Bayes bound in Lemma 3.2. As defined above, for a fixed $w$, the prior distribution $P=\mathcal{N}(0,\sigma^{2}I)$, the posterior distribution $Q$ is the distribution of $\Delta w$ shifted by $w$, that is, $Q=\mathcal{N}(w,\sigma^{2}I)$. Then we have,
@@
$\leq\frac{2}{\sqrt{n}}\left\|H_{l-1}^{*}\right\|_{F}\left\|W_{l}\right\|_{2}\leq 2(B+\epsilon)LM_{1}M_{2}\dfrac{\tau^{l-1}-1}{\tau-1}$
@@
where $\hat{G}=\arg\inf_{G^{\prime}\in\delta_{w}(G)}M\left(f_{w}\left(G^{\prime}\right),i,j\right)$ and $H_{l-1}^{*}$ is the corresponding node representation. Therefore, by the definition, we always have $\hat{L}_{S,\gamma}\left(f_{w}\right)=1$ when $\beta^{l}<\frac{\gamma}{2(B+\epsilon)}$.
Alternatively, if $\beta^{l}>\frac{\gamma\sqrt{m}}{2(B+\epsilon)}$, the term inside the big-O notation in Equation 18 would be
@@
And note that the lower bound in Equation 19 ensures that Equation 17 holds which in turn justifies the applicability of Lemma A.5.
Consider a covering set centered at $C=\left\{\tilde{\beta}_{1},\dots,\tilde{\beta}_{|C|}\right\}$ of the interval in Equation 19 with radius $\frac{1}{l+2}\left(\frac{\gamma}{2(B+\epsilon)}\right)^{1/l}$, where $|C|$ is the size of the set. To make sure that $|\beta-\tilde{\beta}|\leq\frac{1}{l+2}\left(\frac{\gamma}{2(B+\epsilon)}\right)^{1/l}\leq\frac{1}{l+2}\beta$, the minimal size of $C$ should be
@@
$\leq\hat{R}_{S,\gamma}\left(f_{w}\right)$
$\quad\mathcal{O}\left(\sqrt{\frac{(B+\epsilon)^{2}\max\left\{\zeta^{-1},\xi^{1/l}\right\}^{2l+2}l^{2}h\log(lh)\left\|w\right\|_{2}^{2}+\log(m(l+2)/\delta^{\prime})}{\gamma^{2}m}}\right),$
@@
and one of the sufficient conditions is $\frac{l\gamma}{4e(B+\epsilon)(l+1)^{2}}\leq\beta^{2}$. We will see how to satisfy this condition later.
We now compute the KL term in the PAC-Bayes bound in Lemma 3.2. As defined above, for a fixed $w$, the prior distribution $P=\mathcal{N}(0,\sigma^{2}I)$, the posterior distribution $Q$ is the distribution of $\Delta w$ shifted by $w$, that is, $Q=\mathcal{N}(w,\sigma^{2}I)$. Then we have,
@@
where $\hat{G}=\arg\inf_{G^{\prime}\in\delta_{w}(G)}M\left(f_{w}\left(G^{\prime}\right),i,j\right)$ and $H_{l-1}^{*}$ is the corresponding node representation. Therefore, by the definition, we always have $\hat{R}_{S,\gamma}\left(f_{w}\right)=1$ when $\beta^{2}<\frac{\gamma}{2(B+\epsilon)l}$.
Alternatively, if $\beta^{2}>\frac{\gamma\sqrt{m}}{2(B+\epsilon)l}$, the term inside the big-O notation in Equation 21 would be
@@
And note that the lower bound in Equation 22 ensures that Equation 20 holds which in turn justifies the applicability of Lemma A.5.
Consider a covering set $C=\left\{\tilde{\beta}_{1},\dots,\tilde{\beta}_{|C|}\right\}$ of the interval in Equation 22 with radius $\frac{1}{4}\sqrt{\frac{\gamma}{2(B+\epsilon)l}}$, where $|C|$ is the size of the set. To make sure that $|\beta-\tilde{\beta}|\leq\frac{1}{4}\sqrt{\frac{\gamma}{2(B+\epsilon)l}}\leq\frac{1}{4}\beta$, the minimal size of $C$ should be
@@
$\leq\hat{L}_{S,\gamma}\left(f_{w}\right)$
$\quad+\mathcal{O}\left(\sqrt{\frac{(B+\epsilon)^{2}\max\left\{\zeta^{-6},\left(LM_{1}M_{2}\right)^{3}\right\}(l+1)^{4}h\log(lh)\left\|w\right\|_{2}^{2}+\log(m/\delta)}{\gamma^{2}m}}\right).$